Find the average rate of change of x^2+5x from x1=3 and x2=6
thanks!

Found 2 solutions by nerdybill, stanbon:
Answer by nerdybill(7384)

Find the average rate of change of x^2+5x from x1=3 and x2=6
Think of it as the slope between those two points.
f(x) = x^2+5x
f(3) = 3^2+5(3) = 9+15 = 24
f(6) = 6^2+5(6) = 36+30 = 66
.
rise:66-24 = 42
run: 6-3 = 3
.
Average rate of change = 42/3 = 14


Answer by stanbon(75887)

Find the average rate of change of x^2+5x from x1=3 and x2=6
-------------
Average = [f(x1)-f(x2)]/[x1-x2]
-------
Ave = [f(3)-f(6)]/[3-6]
---
Ave = [(3^2+5*3)-(6^2-5*6)]/[-3]
---
Ave = [24-6]/-3
Ave = -6
=================================
Cheers,
Stan H.